# An Attention-Assisted AI Model for Real-Time Underwater Sound Speed Estimation Leveraging Remote Sensing Sea Surface Temperature Data

### v2

@@ III-B Fusion data construction
For a single coordinate, the data fusion structure is shown in figure 2, which includes the real-time SST data, latitude and longitude coordinates, and the first three EOF feature vectors of the historical SSPs. EOF decomposition, also known as principal component analysis, is a statistical method used to analyze structural features in matrix data and extract major data feature quantities. Through EOF decomposition, the change information of the original sound velocity field can be easily condensed into the first few principal components and their corresponding spatial functions, so the main information of the sound velocity field can be expressed only by the first few eigenvectors.
@@
Finally, by fusing data from 8 coordinate points, input training data can be obtained. For location $L_{n-1,m-1}$, the fused data input $\mathcal{F}^{X}_{n-1,m-1}$ can be expressed as:
@@
where $\boldsymbol{T^{S}_{n-1,m-1}}=[T^{s}_{n-1,m-1,1},T^{s}_{n-1,m-1,2},...,T^{s}_{n-1,m-1,H}]^{T}$ is actually a vector composed of H copies of $T^{s}_{n-1,m-1}$ at location $L_{n-1,m-1}$, and $\boldsymbol{L_{n-1,m-1}}$ is similar that $\boldsymbol{L_{n-1,m-1}}=[L_{n-1,m-1,1},L_{n-1,m-1,2},...,L_{n-1,m-1,H}]^{T}$. The label output data is $\mathcal{S}^{Y}_{n,m}=[s^{Y}_{n,m,1},s^{Y}_{n,m,2},...,s^{Y}_{n,m,H}]^{T}$ at location $L_{n,m}$.
@@ III-C SA-MDF-CNN Model
The network structure of SA-MDF-CNN is shown in Figure 3, mainly including attention module and convolutional network module. The first layer is the input layer, the size of which is 1976*6*8, including 6 types of physical quantities (remote sensing SST, latitude, longitude, and the first 3 order eigenvectors) at 8 coordinates with a total depth of 1976 meters. The second layer is a flattening layer, which flattens the three-dimensional fusion data into one-dimensional sequence data as input for the self-attention layer. The third layer comes the self-attention layer, which converts the fusion data into query, key and value vectors through three linear transforms. Then, the dot product of each query vector and each key vector is calculated to obtain the attention score matrix, which is further converted into attention weights through softmax function. Finally, the value vectors are summed using attention weights to obtain the final output features. These layers are designed to enhance the model’s ability of paying attention to the relationship between SSPs in different ocean areas in the input fusion data, so as to better capture the global dependence and key information between fusion data and the sound velocity distribution in the target ocean area.
The fourth and fifth layers are functional layers designed to reshape one-dimensional sequence data into two-dimensional data for subsequent feature extraction through convolution. The convolution layer is designed to further extract the features from the output data of the attention layer, which is better at capturing small-scale relationships of input features. The pooling layer is used to reduce the spatial dimension of the data, thereby reducing the number of parameters and reduce computational complexity, while retaining important feature information. The final two layers are the fully connected layer and the regression layer. The fully connected layer integrates the local features into the global features and outputs them to the regression layer. The RMSE is designed to be the loss function for calculating the gradient of the model parameters and updating weight parameters of the model through back propagation (BP) algorithm:
@@ III-D Multi-head self-attention module
The self-attention mechanism is a special type of attention mechanism used to process sequential data such as text, images, or time series. It allows the model to process the sequence in a way that considers the relationship between each element in the sequence and all other elements. The self-attention mechanism calculates weights by evaluating the similarities between different elements in a sequence to determine how much an element is related to other elements. Specifically, a kind of similarity score is defined to describe the similarity between each element in the sequence and other elements. Then, based on these similarity scores, the weight of each element can be calculated, and a weighted representation can be further obtained by multiplying the weight with the corresponding element. In order to better capture the associations between different types of SSP, multiple attention heads are used in this paper as shown in figure 4.
The diagram presented in Figure 4 illustrates the intricate workings of the multi-head self-attention mechanism, which is a fundamental building block in the architecture of transformer models. Three pivotal components are involved in the self-attention mechanism: the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ vectors. These vectors are derived from the input data through linear transformations, as depicted in the upper right section of the diagram where the input is projected into the query, key, and value spaces using learnable weight matrices.
@@
The mechanism operates by first computing the dot product between the query and each key vector. To prevent the dot products from growing too large and causing numerical instability, especially when the dimensionality of the keys $d_{k}$ is high, each key is normalized by dividing it by $\sqrt{d_{k}}$. This normalization step is shown in the diagram as part of the attention calculation process. Following this, the softmax function is applied to the normalized dot products to generate a set of attention weights. These weights are then used to compute a weighted sum of the value vectors, as represented by the equation:
$\text{Attention}(\mathcal{Q},\mathcal{K},\mathcal{V})=\text{softmax}\left(\frac{\mathcal{Q}\mathcal{K}^{\top}}{\sqrt{d_{k}}}\right)\mathcal{V}$ (10)
This operation is performed for each head in the multi-head self-attention mechanism, as shown in the lower section of the diagram. Each head attends to different parts of the input sequence, allowing the model to capture a diverse range of dependencies and nuances within the data. The outputs from all heads are then concatenated and subjected to a linear transformation, resulting in the final output of the multi-head self-attention mechanism.
@@
Given an input vector $\mathbf{x}\in\mathbb{R}^{dim}$, represented as $\mathcal{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}]$, where $dim$ denotes the number of feature dimensions and $M$ represents the number of sequence samples. Upon receiving a query vector $\mathbf{q}$, the self-attention function is defined as:
$\mathrm{att}(\mathcal{X},\mathbf{q})=\sum_{i=1}^{M}a_{i}\mathbf{x}_{i}=\sum_{i=1}^{M}\frac{\exp(\mathrm{s}(\mathbf{k}_{i},\mathbf{q}))}{\sum_{j=1}^{M}\exp(\mathrm{s}(\mathbf{k}_{j},\mathbf{q}))}\mathbf{v}_{i},$ (11)
@@
where $a_{i}$ is the $i$th attention weight, computed by the softmax function over the scaled dot product of the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$. The function $\mathrm{s}(\mathbf{k}_{i},\mathbf{q})$ represents the similarity score between the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$, which is typically calculated as the dot product $\mathbf{k}_{i}^{\top}\mathbf{q}$ divided by the square root of the key vector’s dimension $\sqrt{d_{k}}$. The value vector $\mathbf{v}_{i}$ corresponds to the $i$th sample in the input sequence, and the output of the self-attention function is a weighted sum of these value vectors, where the weights are determined by the attention mechanism.
@@
In a standard self-attention module, the multi-head attention mechanism employs multiple query vectors $\mathcal{Q}$ to select multiple sets of information from the input data in parallel. Specifically, each head independently computes the attention weights between the query vectors $\mathcal{Q}$, key vectors $\mathcal{K}$, and value vectors $\mathcal{V}$ to extract different subsets of information. The output of the multi-head self-attention module is calculated by concatenating the outputs of all heads and then applying a linear transformation. The mathematical expression is:
@@
where $\mathrm{head_{i}}=\text{Attention}(\mathcal{QW}_{i}^{Q},\mathcal{KW}_{i}^{K},\mathcal{VW}_{i}^{V})$, $\mathcal{W}_{i}^{Q}\in\mathbb{R}^{d_{k}},\mathcal{W}_{i}^{K}\in\mathbb{R}^{d_{k}},\mathcal{W}_{i}^{V}\in\mathbb{R}^{d_{v}}$, and $\mathcal{W}^{o}\in\mathbb{R}^{hd_{v}}$ is a projection of the parameter matrix.
@@ IV Results and discussions
To test the effectiveness of the proposed SA-MDF-CNN model, the monthly average fusion data from July 2015 to 2019 were selected as input training data and the monthly average SSP from 2020 were set to be the output label data. Experimental results were compared with other widely used methods for constructing spatial sound velocity distribution, including convolutional neural network (CNN), spatial interpolation (SITP) and mean values (MEAN), respectively. The model parameter settings are given in Table I.
@@ IV-A Accuracy Performance
To test the accuracy performance of proposed model in estimating SSPs, real-time remote sensing SST data from different locations and depths were fused with EOF feature vectors to construct the sound velocity field. The remote sensing SST data are shown in figure 5, and a visual display of the fusion data is given in figure 6.
@@
Figure 7: Comparison of real-time estimated SSP results of different algorithms with depth of 1976 m, where (a) is the SA-MDF-CNN algorithm, (b) is the CNN algorithm, (c) is the spatial interpolation algorithm, and (d) is the mean value method.
@@
The real-time SSP estimation results of SA-MDF-CNN algorithm with depths ranging from 0 to 1976 meters within $7^{\circ}N$ to $28^{\circ}N$, $150^{\circ}E$ to $165^{\circ}E$ are compared with those of CNN, SITP and mean values method as shown in figure 7. From the intuitive representation of the sound velocity heat-map, it can be seen that the sound velocity distribution estimated by SA-MDF-CNN and CNN has a significant improvement in accuracy compared to SITP and Mean value methods, and the estimated values of the proposed model are more similar to the sound velocity distribution of real samples. A comprehensive evaluation of the accuracy performance of various algorithms across different locations at depths ranging from 0 to 1976 meters is presented in Table II. Intuitive observation shows that the average RMSE results of SA-MDF-CNN, CNN, SITP, and Mean value methods are 0.1468, 0.1661, 0.2326, and 0.2447, respectively. Notably, the RMSE of SA-MDF-CNN for real-time SSP estimation is approximately 12% lower than that of CNN and approximately 30-40% lower than that of traditional SITP and mean value methods. Figure 8 gives a comparison of sound velocity estimation curves at $7.5^{\circ}N,156.5^{\circ}E$. In figure 8 (a), the curve of SA-MDF-CNN is the closest to the real SSP, and in figure 8 (b), the error mainly manifests in the shallow ocean area within 500 meters because the sound velocity in shallow water is more severely and irregularly affected by temperature. However, among all methods the error disturbance of SA-MDF-CNN is the smallest.
@@
To further test the accuracy and performance of the model in estimating the sound velocity distribution in shallow water areas, the real-time SSP estimation results of SA-MDF-CNN algorithm at 200 m, 300 m and 500 m depths are compared with those of CNN, SITP and mean values method as shown in figure 9. As can be seen from the enlarged part of each sub-figure, the estimated SSP curves by SA-MDF-CNN at 200 m, 300 m and 500 m depths are significantly closer to the real SSP prediction results than those of CNN, SITP and mean method at the same ocean area. Specifically, an in-depth investigation into the performance of different algorithms in shallow ocean areas at varying depths and locations was conducted, with the real-time SSP estimation RMSE results compared and displayed in Table LABEL:table3. The largest RMSE is observed at a depth of 200 meters, which can be attributed to the increased interference from real-time sea conditions in the shallow sea environment, such as wind speed, wave height, tsunami, and other factors. Concurrently, the experimental results indicate that the estimation accuracy of the proposed SA-MDF-CNN model at 200 meters depth is enhanced by approximately 13% compared to CNN and by approximately 53% compared to SITP and mean value method. At an ocean depth of 300 meters, the proposed algorithm demonstrates an improvement of about 14% over CNN and about 42% over SITP and mean value method. At 500 meters depth, the proposed model shows an improvement of about 14% compared to CNN and about 49% compared to SITP and mean value method. These experimental results collectively illustrate that the proposed model maintains significant precision advantages even in shallow sea environments of varying depths, thereby underscoring its applicability and effectiveness. The results indicate that SA-MDF-CNN is also applicable even in shallow ocean environment with different depths, and still has smaller RMSE compared with other algorithms, which means that the model still has good feature capture ability in shallow ocean scenes with significant changes in sound velocity. The average absolute error distribution comparison of real-time estimated SSP of different algorithms tested in the set of 200m and 500m depths is given in Figure 10. The error fluctuation of SA-MDF-CNN is the most stable, which indicates that SA-MDF-CNN still has better robust performance under shallow sea conditions with different depths.

### v3

@@ III-B Fusion data construction
For a specific single coordinate, the configuration of data fusion is presented in figure 2. This configuration encompasses real-time SST data, coordinates of latitude and longitude, along with the principal three eigen-vectors of the EOF of historical SSPs. EOF decomposition, alternatively referred to as principal component analysis, is a statistical approach employed for examining the structural characteristics within matrix data and extracting the primary quantitative features of the data. By means of EOF decomposition, the variation information of the original sound velocity field can be readily compressed into a few principal components and their corresponding spatial functions. As a result, the essential information of the sound velocity field can be represented solely by the primary few eigen-vectors.
@@
Finally, by fusing data from 8 coordinates, input training data can be obtained. For location $L_{n-1,m-1}$, the fused data input $\mathcal{F}^{X}_{n-1,m-1}$ can be expressed as:
@@
where $\boldsymbol{T^{S}_{n-1,m-1}}=[T^{s}_{n-1,m-1,1},...,T^{s}_{n-1,m-1,H}]^{T}$ is actually a vector composed of H copies of $T^{s}_{n-1,m-1}$ at location $L_{n-1,m-1}$, and $\boldsymbol{L_{n-1,m-1}}$ is similar that $\boldsymbol{L_{n-1,m-1}}=[L_{n-1,m-1,1},L_{n-1,m-1,2},...,L_{n-1,m-1,H}]^{T}$. The label output data is $\mathcal{S}^{Y}_{n,m}=[s^{Y}_{n,m,1},s^{Y}_{n,m,2},...,s^{Y}_{n,m,H}]^{T}$ at location $L_{n,m}$.
@@ III-C SA-MDF-CNN Model
The architectural framework of the SA-MDF-CNN is illustrated in Figure 3, mainly including attention module and convolutional network module. The initial layer is the input layer, the size of which is 1976*6*8, including 6 types of physical quantities (remote sensing SST, latitude, longitude, and the first 3 order eigenvectors) at 8 coordinates with a total depth of 1976 meters. Then the flattening layer serves as a flattening operation, transmuting three-dimensional fusion data into one-dimensional sequence data as input for the self-attention mechanism. The following module is designated as the self-attention layer, which transmutes integrated data into query, key, and value vectors via three distinct linear transformation processes. Subsequently, the inner product between every query vector and each key vector is computed, resulting in an attention score matrix. This matrix is subsequently transformed into attention weights via the softmax function. The final output features are generated through the aggregation of value vectors, facilitated by the application of attention weights. The architectural design of these layers is intended to augment the model’s capacity for focusing on the inter - relationships between SSPs situated across disparate oceanic zones within the input fusion data, so as to better capture the global dependence and key information between fusion data and the sound velocity distribution in the target ocean area.
The fourth and fifth layers are functional layers designed to reshape one-dimensional sequence data into two-dimensional data for subsequent feature extraction through convolution. The convolution layer is intended to derive additional features from the resultant data of the attention layer, which is better at capturing small-scale relationships of input features. The pooling layer serves to diminish the spatial dimensionality, thereby reducing the number of parameters and reduce computational complexity, while retaining important feature information. The penultimate and ultimate layers comprise the fully connected and the regression layer. The fully connected layer integrates local features into global representations, subsequently passing them to the regression layer. The RMSE is designated as the loss function, which is employed to compute the gradient of the model parameters and update the model’s weight parameters via the back propagation algorithm.
@@ III-D Multi-head self-attention module
The self-attention mechanism constitutes a distinct variant of the attention mechanism, primarily employed for processing temporal data such as text, images, or time-series data. It enables the model to analyze the sequence in a manner that takes into account the inter-element relationships within the sequence. The self-attention mechanism calculates weights by evaluating the similarities between different elements in a sequence to determine how much an element is related to other elements. Specifically, a kind of similarity score is defined to describe the similarity between each element in the sequence and other elements. Then, based on these similarity scores, the weight of each element can be calculated, and a weighted representation can be further obtained by multiplying the weight with the corresponding element. In order to better capture the associations between different types of SSP, multiple attention heads are used in this paper as shown in figure 4.
The diagram presented in Figure 4 illustrates the intricate workings of the multi-head self-attention mechanism, which is a fundamental building block in the architecture of transformer models, depending on three integral elements: the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ vectors. These vectors are derived from the input data through linear transformations, as depicted in the upper right section of the diagram.
@@
The operation of the mechanism commences with the calculation of the dot-product between the query and every key vector. To avoid an excessive increase in the dot-products that could trigger numerical instability, particularly when the dimensionality of the keys $d_{k}$ is substantial, each key is standardized through division by $\sqrt{d_{k}}$. This standardization stage is depicted in the diagram as an element of the attention computation procedure. Afterward, the softmax function is employed on the standardized dot-products to yield a collection of attention weights. Subsequently, these weights are utilized to calculate a weighted sum of the value vectors, as expressed by the equation:
$\text{Att}(\mathcal{Q},\mathcal{K},\mathcal{V})=\text{softmax}\left(\frac{\mathcal{Q}\mathcal{K}^{\top}}{\sqrt{d_{k}}}\right)\mathcal{V}$ (10)
This operation is carried out for every head within the attention mechanism, as illustrated in the lower part of the diagram. Each head focuses on distinct portions of the input sequence, enabling the model to grasp a wide variety of dependencies and subtleties in the data. Subsequently, the outputs of all heads are concatenated and undergo a linear transformation, culminating in the ultimate output of the attention mechanism.
@@
Given an input vector $\mathbf{x}\in\mathbb{R}^{dim}$, represented as $\mathcal{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}]$, where $dim$ signifies the number of feature dimensions and $M$ indicates the number of sequence samples. Once a query vector $\mathbf{q}$ is received, the self-attention function is formulated as:
$\mathrm{Sa}(\mathcal{X},\mathbf{q})=\sum_{i=1}^{M}a_{i}\mathbf{x}_{i}=\sum_{i=1}^{M}\frac{\exp(\mathrm{s}(\mathbf{k}_{i},\mathbf{q}))}{\sum_{j=1}^{M}\exp(\mathrm{s}(\mathbf{k}_{j},\mathbf{q}))}\mathbf{v}_{i},$ (11)
@@
where $a_{i}$ is the $i$th attention weight, computed by the softmax function over the scaled dot-product of the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$. The function $\mathrm{s}(\mathbf{k}_{i},\mathbf{q})$ represents the similarity score between the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$, which is typically calculated as the dot-product $\mathbf{k}_{i}^{\top}\mathbf{q}$ divided by the square root of the key vector’s dimension $\sqrt{d_{k}}$. The value vector $\mathbf{v}_{i}$ corresponds to the $i$th sample in the input sequence.
@@
In the self-attention module, the multi-head attention mechanism employs multiple query vectors $\mathcal{Q}$ to select multiple sets of information from the input data in parallel. Specifically, each head independently computes the attention weights between the query vectors $\mathcal{Q}$, key vectors $\mathcal{K}$, and value vectors $\mathcal{V}$ to extract different subsets of information. The resultant representation of the multi-head self-attention mechanism is derived through the concatenation of results from each attention head, followed by a linear projection operation. The mathematical expression is:
@@
where $\mathrm{head_{i}}=\text{Att}(\mathcal{QW}_{i}^{Q},\mathcal{KW}_{i}^{K},\mathcal{VW}_{i}^{V})$, $\mathcal{W}_{i}^{Q}\in\mathbb{R}^{d_{k}}$, $\mathcal{W}_{i}^{K}\in\mathbb{R}^{d_{k}}$, $\mathcal{W}_{i}^{V}\in\mathbb{R}^{d_{v}}$, and $\mathcal{W}^{o}\in\mathbb{R}^{hd_{v}}$ is a projection of the parameter matrix.
@@ IV Results and discussions
To test the effectiveness of the proposed SA-MDF-CNN model, the monthly average fusion data from July 2015 to 2019 were selected as input training data and the monthly average SSPs from 2020 were set to be the output label data. Experimental results were compared with other widely used methods for constructing spatial sound velocity distribution, including convolutional neural network (CNN), spatial interpolation (SITP) and mean values (MEAN), respectively. The model parameter settings are given in Table I.
@@ IV-A Accuracy Performance
To generate the fusion data necessary for evaluating the accuracy of the recommended model in estimating the sound velocity field, real-time remote sensing SST data from multiple locations were combined with the EOF feature vector to reconstruct the sound velocity field. Figure 5 shows the remote sensing SST data, and figure 6 presents a visual depiction of the fused data.
@@
Figure 7: Evaluation of real-time SSP estimation outcomes across diverse algorithmic approaches at 1976 meter bathymetric measurement, where (a) is the SA-MDF-CNN algorithm, (b) is the CNN algorithm, (c) is the spatial interpolation algorithm, and (d) is the mean value method.
@@
The real-time SSP estimation results of SA-MDF-CNN algorithm with depths ranging from 0 to 1976 meters within $7^{\circ}N$ to $28^{\circ}N$, $150^{\circ}E$ to $165^{\circ}E$ are compared with those of CNN, SITP and mean values method as shown in figure 7. From the intuitive representation of the sound velocity heat-map, the sound velocity distribution estimated by the SA-MDF-CNN and CNN models exhibits a substantial enhancement in accuracy relative to the SITP and Mean value methods as evidenced by the results, and the estimated values of the proposed model are more similar to the sound velocity distribution of real samples. A comprehensive evaluation of the accuracy performance of various algorithms across different locations at depths ranging from 0 to 1976 meters is presented in Table II. Intuitive observation shows that the average RMSE results of SA-MDF-CNN, CNN, SITP, and mean value methods are 0.1468, 0.1661, 0.2326, and 0.2447, respectively. Notably, the RMSE of SA-MDF-CNN for real-time SSP estimation is approximately 12% lower than that of CNN and approximately 30-40% lower than that of traditional SITP and mean value methods. Figure 8 gives a comparison of sound velocity estimation curves at $7.5^{\circ}N,156.5^{\circ}E$. In figure 8 (a), the curve of SA-MDF-CNN is the closest to the real SSP, and in figure 8 (b), the error mainly manifests in the shallow ocean area within 500 meters because the sound velocity in shallow water is more severely and irregularly affected by temperature. However, among all methods the error disturbance of SA-MDF-CNN is the smallest.
@@
To rigorously evaluate the precision and performance of the model for reconstructing acoustic velocity profiles in coastal marine environments, a comparative analysis was conducted between the instantaneous SSP reconstructions generated by the SA-MDF-CNN framework at 200 m, 300 m, and 500 m depth layers and those produced by conventional CNN architectures, the SITP technique, and statistical averaging approaches, as illustrated in Figure 9. The magnified sections of individual subplots reveal that SA-MDF-CNN-derived SSP estimations at 200 m, 300 m, and 500 m depth layers demonstrate enhanced alignment with ground-truth SSP predictions when compared to conventional CNN architectures, SITP algorithms, and mean method within identical marine regions. Specifically, an in-depth investigation into the efficacy of various algorithms within shallow ocean areas at varying depths and locations was conducted, with the real-time SSP estimation RMSE results compared and displayed in Table LABEL:table3. The largest RMSE is observed at a depth of 200 meters, which can be attributed to the increased interference from real-time sea conditions in the shallow sea environment, such as wind speed, wave height, tsunami, and other factors. Concurrently, the experimental results indicate that the estimation accuracy of the proposed SA-MDF-CNN model at 200 meters depth is enhanced by approximately 13% compared to CNN and by approximately 53% compared to SITP and mean value method. At an ocean depth of 300 meters, the proposed algorithm demonstrates an improvement of about 14% over CNN and about 42% over SITP and mean value method. At 500 meters depth, the proposed model shows an improvement of about 14% compared to CNN and about 49% compared to SITP and mean value method. These experimental results collectively illustrate that the proposed model maintains significant precision advantages even in shallow sea environments of varying depths, thereby underscoring its applicability and effectiveness. Experimental findings reveal the SA-MDF-CNN architecture maintains operational efficacy across heterogeneous depth profiles within coastal marine conditions, while exhibiting RMSE relative to conventional methodologies, which means that the model still has good feature capture ability in shallow ocean scenes with significant changes in sound velocity. The average absolute error distribution comparison of real-time estimated SSP of different algorithms tested in the set of 200m and 500m depths is given in Figure 10. The error fluctuation of SA-MDF-CNN is the most stable, which indicates that SA-MDF-CNN still has better robust performance under shallow sea conditions with different depths.